What volume of a 0.126 M potassium hydroxide solution is required to neutralize 22.0 mL of a 0.139 M nitric acid solution?


___________mL potassium hydroxide

1
Expert's answer
2021-10-11T06:10:00-0400

v=0.139/22=0.006mol

V=0.126/0.006=21ml